# Pair Production in time-dependent Electric field at Finite times

### v4

@@ Abstract
Particle Production, Dynamical Schwinger effect, Strong-Field QED
@@ I Introduction
The concept of pair production in an electromagnetic field has its roots in the mid-1920s after the invention of quantum mechanics, with the formulation of the relativistic wave equation for electrons by Paul Dirac in 1928 Dirac (1928). The Dirac Sea model was proposed to explain the enigma of negative energy solutions. F. Sauter’s work in 1931 demonstrated that strong electric fields can lead to pair creation through tunneling with exponential suppression Sauter (1931). This paved the way for quantum field theory, recognizing the vacuum as a polarizable medium influenced by constant fluctuations. In 1935, W. Heisenberg and H. Euler further explored the peculiarities of the Dirac equation, revealing non-linear modifications in Maxwell’s equations due to the interaction of electromagnetic fields with the electron vacuum loopHeisenberg and Euler (1936). J. Schwinger’s groundbreaking work in 1951 precisely calculated the imaginary part of the one-loop effective Lagrangian in the presence of a static electric field Schwinger (1951). As a result of his seminal work, the phenomenon of vacuum pair creation by electric fields has since become widely known as the Schwinger effect, and it is also famously referred to as the Sauter-Schwinger effect in recognition of F. Sauter’s prior work on solving the Dirac equation in the presence of an electric field Sauter (1932). Schwinger’s pioneering calculation opened up new avenues of research in quantum field theory and has profoundly impacted our understanding of particle physics in the presence of strong fields. This extraordinary property of quantum vacuum producing spontaneous particle-antiparticle pairs has far-reaching implications for understanding the generation of particle-antiparticle pairs in the presence of a strong electric field Greiner et al. (1985); particle creation in the expanding universeParker (1968); black hole evaporation as a result of Hawking radiation Hawking (1975); Parikh and Wilczek (2000); and Unruh radiation, in which particle production is seen by an accelerating observerUnruh (1976); Kim (2016). The study of generating electron-positron pairs through a spatially constant electrical background field was extended to the electric field with various time dependences. In the 1970s, researchers explored the occurrence of pair production from the vacuum in the presence of an oscillating time-dependent electric field Popov (1971); Marinov and Popov (1977). Their investigations revealed different qualitative behaviors for this process under various interaction regimes. The interaction regimes can be distinguished by the value of the dimensionless Keldysh parameter $\gamma=\frac{m\omega}{|e|E_{0}}$ Keldysh (1965), with field amplitude $E_{0}$, field frequency $\omega$, electron charge $e$, and mass $m$. When $\gamma>>1$, the process probability shows a perturbative power-law scaling with field intensity. Instead, for $\gamma<<1$, it exhibits a manifestly non-perturbative exponential dependence on $\frac{1}{E_{0}}$, similar to the case of a constant electric field Nikishov (1969). Particle production in a spatially homogeneous single pulse of an electric field has been explored Narozhnyi and Nikishov (1970); Nikishov (2001), and methods for tackling particle creation in an arbitrary time-dependent electric field have been developed Narozhnyi and Nikishov (1970); Brezin and Itzykson (1970); Bagrov et al. (1975). Because a single pulse of an electric field is an idealized form of an electric field created by two colliding laser beams, particle generation in an alternating electric field has been explored for a more realistic scenario Mostepanenko and Frolov (1974); Narozhnyi and Nikishov (1973). However, it was found that the mass of the created particle, via the Schwinger mechanism, exponentially suppresses the pair generation rate, necessitating a very strong electric field to observe this phenomenon. Therefore, this event has not yet been the subject of any experimental observations. This makes it unclear how well theoretical prediction captures the physics of pair production. Nevertheless, the study of pair production in strong electric fields subject has attracted sustained interest from theoreticians in recent years due to the extraordinary progress in the development of the ultra-intense lasers technique and the strong field QED experimental studies that are planned at upcoming high-intensity laser facilities, such as the European X-Ray Free-Electron Laser Andruszkow et al. (2000), the Extreme-Light Infrastructure Heinzl and Ilderton (2009); Dunne (2009), the Exawatt Center for Extreme Light Studies Khazanov and et al. (2023). The electric field strength is fast getting closer to the critical value $E_{\text{c}}$. Additionally, it is suggested that the Schwinger mechanism can be indirectly tested in the condensed matter system of a single monolayer of graphene, where the electrons are roughly characterized by the massless pseudo-relativistic Dirac equation Allor et al. (2008); Klimchitskaya and Mostepanenko (2013); Fillion-Gourdeau and MacLean (2015).
@@
Particle production can be viewed as evolving a quantum system from an initial equilibrium configuration to a new final equilibrium configuration via an intermediate non-equilibrium evolution caused by a strong field background. In intermediate non-equilibrium states, when matter fields interact with a time-dependent external field, the classical Hamiltonian loses its time-translation invariance, leading to the evolution of the vacuum state into a state where mixing between positive and negative energy states occurs. Special care is needed in identifying “positive” and “negative” energy states in the presence of an external electric field. Thus, energy is no longer a good quantum number, and there is no well-defined quantum number to characterize a “particle”. However, the standard way of describing particle creation by an external field involves the asymptotic analysis of particle states in the remote past (or in-states, before the external field is switched on) and the distant future (or out-states, long after the interaction with the external field has ended). This method is well understood using the quantum field operator expression in terms of creation and annihilation operators, which are connected to single-particle states in both the past and the future. Then, by relating the two sets of operators in the past and future, one can derive an equation for the S-matrix of the process. This equation, in turn, determines the number of particles produced in the process. Various methodologies have been developed to investigate pair production in the presence of strong fields. These include the proper-time method Schwinger (1951); DeWitt (1975), the canonical method Mottola (1985); Tomaras et al. (2001, 2000), Green’s function techniques Gavrilov and Gitman (1996), semiclassical tunneling Parikh and Wilczek (2000), the Schrödinger-functional approach Hallin and Liljenberg (1995), functional techniques Fried and Woodard (2002); Avan et al. (2003); Ilderton (2014),the Wigner formalismZhuang and Heinz (1996); Hebenstreit et al. (2011a, b),mean-field treatments Smolyansky et al. (1998), and worldline instanton methods Affleck et al. (1982); Kim and Page (2002). Studies on pair production in intense laser fields Alkofer et al. (2001); Roberts et al. (2002) have employed analytical and numerical approaches to predict pair-production rates. However, these investigations primarily focus on pair formation at asymptotic times, which alone is insufficient for fully capturing the dynamical aspects of pair production and its time-dependent effects. Several authors have derived expressions for the number of pairs created after a time $T$, where $T$ extends beyond the duration of the electric field pulse or the point at which the field vanishes Gavrilov and Gitman (1996); Di Piazza (2004); Melike et al. (2012); Ilderton (2022); Adorno et al. (2017). These expressions provide a reliable approximation of the actual pair count. Understanding pair production dynamics from the initial to final states requires a framework that accounts for transient and non-equilibrium effects. In this context, the distribution function of quasiparticles, governed by the quantum kinetic equation (QKE), offers an effective description of the time-dependent evolution of pair production Blaschke et al. (2015). Defining the particle distribution function in the quasiparticle basis is both physically meaningful and intuitive. This approach is grounded in the straightforward physics of how a time-dependent field distorts the time invariance symmetry of the Hamiltonian, leading to mixing between momentum states of positive and negative energies. This method is particularly relevant for exploring the complex dynamical behavior of pair production in spatially uniform, time-dependent electric fields. Specifically, it has been demonstrated that the temporal evolution of the quasiparticle distribution function in momentum space provides insights into the asymptotic states of the quantum field and offers a comprehensive description of the process during the interaction between the matter field and the strong field background Blaschke et al. (2017). It was shown that the distribution function evolves through three distinct temporal stages: the quasi-electron-positron plasma (QEPP) stage, the transient stage, and finally the residual electron-positron plasma (REPP) stage Banerjee and Singh (2022, 2019). In examining the pair production process, both the QEPP stage, where quasi-particles are formed, and the REPP stage, where real particles emerge, are crucial. While the Schwinger mechanism is primarily known for producing elementary particles in a vacuum under external fields, it can also generate quasiparticle excitations in exotic materials such as graphene Allor et al. (2008); Zubkov (2012); Gavrilov et al. (2012); Klimchitskaya and Mostepanenko (2013); Fillion-Gourdeau and MacLean (2015), which are experimentally more accessible, and in superconductors Solinas et al. (2021). Moreover, while ultracold atoms in optical lattices have been investigated as simulators for the Schwinger mechanism Kasper et al. (2016); Halimeh et al. (2025), they offer an alternative platform to study these phenomena Schützhold (2025).
@@
In the present work, we consider the production of electron-positron pairs from the vacuum in a time-varying, spatially uniform pulsed electric field given by $E\left(t\right)=E_{0}\textrm{sech}^{2}\left(\frac{t}{\tau}\right)$, with amplitude $E_{0}$ and pulse duration $\tau$. Such background field has received extensive attention in the literature Narozhnyi and Nikishov (1970); Dunne and Hall (1998); Balantekin et al. (1991); Gavrilov and Gitman (1996); Nikishov (2004); Levai and Skokov (2010); Klimchitskaya and Mostepanenko (2013), with a focus on the asymptotic behavior of the probability of pair production. To this end, we study the evolution of the single-particle distribution function, $f(\bm{p},t)$, which is rigorously derived by canonical quantization of the Dirac field and subsequent Bogoliubov transformation to a quasi-particle representation Smolyansky et al. (2017a). The distribution function exhibits three distinct dynamical stagesBanerjee and Singh (2022); Otto et al. (2016); Nousch et al. (2017); Smolyansky et al. (2017a). By analyzing the temporal evolution of $f(\bm{p},t)$, we find that the emergence and characteristics of these dynamical stages—QEPP, transient, and REPP—are qualitatively and quantitatively influenced by both longitudinal and transverse momentum values. Next, we analyze the longitudinal momentum spectrum (LMS) of the particles at finite times. In the tunneling regime ($\gamma<1$), we observe oscillatory behavior in the LMS at time $t>2\tau$, and this oscillation pattern continuously changes from $t>2\tau$ up to $t<6\tau$ for $\gamma=0.5$. This oscillatory behavior at finite times clearly illustrates the quantum interference effects arising from various dynamical processes and channels contributing to the creation of (quasi-)particles with a given momentum. It can also be understood in terms of the vacuum polarization function $u(\bm{p},t)$ and its counterpart, the depolarization function $v(\bm{p},t)$, which encapsulate the underlying particle production dynamics. We emphasize that the oscillations seen in the LMS are not artifacts but rather possess significant physical relevance. In the multi-photon regime, we observe that LMS at finite times near $t=3\tau$ exhibits a multi-modal structure. Furthermore, we develop an analytical framework valid for finite times $t>\tau$. Specifically, we derive an analytical expression for the single-particle distribution function as a power series in the small parameter $(1-y)$, where $y=\frac{1}{2}(1+\tanh{(\frac{t}{\tau})})$. The intriguing dynamical features of the momentum distribution function at finite times arise from a function appearing in this expansion. The finite-time behavior of the momentum spectra can be understood in terms of functions that reveal three distinct phases, which are discussed in detail in the results section. Additionally, we explored how the transverse momentum affects LMS and discovered that the oscillatory behavior depends on the value of $p_{\perp}$ at that time. Subsequently, we analyzed the temporal progression of transverse momentum spectrum(TMS) of generated pairs, providing valuable insights into the pair production process. At finite times, the transverse momentum spectra reaches its peak at a nonzero value, reflecting the influence of multiphoton transitions in quasiparticle creation.
@@
The preliminary results of this work were presented in Refs. Sah and Singh (2023, 2024). After seeing the report in Ref. Sah and Singh (2023), one of the authors of Ref. Diez et al. (2023), Kohlfürst, informed us in a private communication about their findings, which exhibit some overlap with the results presented here. In this manuscript, we compare and discuss our findings in relation to those of Ref. Diez et al. (2023) whenever applicable.
@@
Throughout the paper, we use natural units and set $\hslash=c=m=|e|=1$, the electric charge $e<0$, and express all variables in terms of the electron mass unit.
@@ II Theory
In this section, we review the canonical quantization method within the context of a strong electric field, as discussed in Grib et al. (1972). This approach offers valuable insights into pair production phenomena, including detailed information on particle number and momentum distribution at each moment. To begin, we establish solutions for the single-particle scenario governed by the Dirac equation in the presence of such a field. We start by writing the Dirac equation for a particle in an electromagnetic field, which takes the following form:
@@
where $A^{\mu}$ is the four-vector potential, $\Psi(\bm{x},t)$ is the four-component spinor. For $\gamma$-matrices we chose the weyl basis Ryder (1996)
$\gamma^{0}=\begin{pmatrix}\mathbb{I}&0\\ 0&-\mathbb{I}\end{pmatrix},\gamma^{i}=\begin{pmatrix}0&-\sigma^{i}\\ \sigma^{i}&0\end{pmatrix},i=1,2,3,$ (2)
where $\mathbb{I}$ is the identity matrix and $\sigma^{i}$ are the Pauli matrices. The $\gamma$ matrices satisfy the anti-commutation relations:
@@
$[(i\partial_{\mu}-eA_{\mu})^{2}-\frac{e}{2}\sigma^{\mu\nu}\mathcal{F}_{\mu\nu}-m^{2}]\chi(\bm{x},t)=0,$ (6)
where $\mathcal{F}_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$ is the field strength tensor, $\sigma^{\mu\nu}=\frac{i}{2}[\gamma^{\mu},\gamma^{\nu}]$ and $\chi(\bm{x},t)$ is a four component spinor.
@@
Spatial homogeneity implies that the solutions can be expressed in the form of
@@
Subsequently, Eq. (7) becomes
$\Bigl{(}\partial_{t}^{2}+\textrm{i}eE(t)\gamma^{0}\gamma^{3}+\omega^{2}(\bm{p},t)\Bigr{)}\chi_{\bm{p}}(t)=0,$ (9)
where
$\omega^{2}(\bm{p},t)=p_{\perp}^{2}+(p_{\parallel}-eA(t))^{2}+m^{2}$ (10)
is expressed in terms of the transverse energy $\epsilon^{2}_{\perp}(p_{\perp})=m^{2}+p_{\perp}^{2}$,longitudinal $p_{\|}=p_{3}$ is the momentum component to the field direction and $p_{\perp}=\sqrt{p_{1}^{2}+p_{2}^{2}}$ is the modulus of the vector ${\bm{p}}_{\perp}$ perpendicular to the field.
@@
These solutions will be interpreted as describing an electron ($\lambda=+$) and its anti-particle, i.e., a positron ($\lambda=-$).
@@
where the spinor solutions, $\Psi_{{\bm{p}}r}^{(\lambda)}(\bm{x},t)$ are normalised according to the product:
@@
Now, the Hamiltonian can be calculated from the energy-momentum tensor which yields,
@@
In order to calculate the spectrum, the Hamiltonian is diagonalized through a basis transformation to the quasi-particle representation using the new time-dependent operators $\hat{B}_{{\bm{p}}r}(t)$ and $\hat{D}_{{\bm{p}}r}(t)$. The relation between the $\hat{b}_{{\bm{p}}r},\hat{d}_{{\bm{p}}r}$and $\hat{B}_{{\bm{p}}r}(t),\hat{D}_{{\bm{p}}r}(t)$ operators is given by the Bogoliubov transformation Grib et al. (1972); Schmidt et al. (1998)
@@
$\hat{D}_{{\bm{p}}r}(t)$ $=\alpha_{-{\bm{p}}}(t)\hat{d}_{{\bm{p}}r}-\beta_{-{\bm{p}}}(t)\hat{b}_{-{\bm{p}}r}^{\dagger}.$ (29)
@@
Here, the operators $\hat{B}_{{\bm{p}}r}(t)$ and $\hat{D}_{{\bm{p}}r}(t)$ describe the creation and annihilation of quasiparticles at time $t$ with respect to the instantaneous vacuum state. The substitution of Eqs. (28) and (29) into Eq. (21) results in a new representation of the field operator
@@
$\phi_{\bm{p}}^{(\lambda)}(t)=\frac{e^{-\textrm{i}\lambda\Theta_{\bm{p}}(t)}}{\sqrt{2\omega(\bm{p},t)\bigl{(}\omega(\bm{p},t)-\lambda P(p_{\parallel},t)\bigr{)}}},$ (35)
@@
We now define the occupation number of electrons in the given eigenmode $\bm{p}r$ of the fermionic field using the time-dependent creation and annihilation operators for the initial vacuum state is defined as
$f_{r}(\bm{p},t)$ $=\langle{0_{in}}|\hat{B}_{{\bm{p}}r}^{\dagger}(t)\hat{B}_{{\bm{p}}r}(t)|{0_{in}}\rangle$ (40)
@@
$\bar{f}_{r}(-\bm{p},t)$ $=\langle{0_{in}}|\hat{D}_{-{\bm{p}}r}^{\dagger}(t)\hat{D}_{-{\bm{p}}r}(t)|{0_{in}}\rangle$ (41)

### v5

@@ I Introduction
The concept of pair production in an electromagnetic field has its roots in the mid-1920s after the invention of quantum mechanics, with the formulation of the relativistic wave equation for electrons by Paul Dirac in 1928 Dirac (1928). The Dirac Sea model was proposed to explain the enigma of negative energy solutions. F. Sauter’s work in 1931 demonstrated that strong electric fields can lead to pair creation through tunneling with exponential suppression Sauter (1931). This paved the way for quantum field theory, recognizing the vacuum as a polarizable medium influenced by constant fluctuations. In 1935, W. Heisenberg and H. Euler further explored the peculiarities of the Dirac equation, revealing non-linear modifications in Maxwell’s equations due to the interaction of electromagnetic fields with the electron vacuum loopHeisenberg and Euler (1936). J. Schwinger’s groundbreaking work in 1951 precisely calculated the imaginary part of the one-loop effective Lagrangian in the presence of a static electric field Schwinger (1951). As a result of his seminal work, the phenomenon of vacuum pair creation by electric fields has since become widely known as the Schwinger effect, and it is also famously referred to as the Sauter-Schwinger effect in recognition of F. Sauter’s prior work on solving the Dirac equation in the presence of an electric field Sauter (1932).
Schwinger’s pioneering calculation opened up new avenues of research in quantum field theory and has profoundly impacted our understanding of particle physics in the presence of strong fields. This extraordinary property of quantum vacuum producing spontaneous particle-antiparticle pairs has far-reaching implications for understanding the generation of particle-antiparticle pairs in the presence of a strong electric field Greiner et al. (1985); particle creation in the expanding universeParker (1968); black hole evaporation as a result of Hawking radiation Hawking (1975); Parikh and Wilczek (2000); and Unruh radiation, in which particle production is seen by an accelerating observerUnruh (1976); Kim (2016). The study of generating electron-positron pairs through a spatially constant electrical background field was extended to the electric field with various time dependences. In the 1970s, researchers explored the occurrence of pair production from the vacuum in the presence of an oscillating time-dependent electric field Popov (1971); Marinov and Popov (1977). Their investigations revealed different qualitative behaviors for this process under various interaction regimes. The interaction regimes can be distinguished by the value of the dimensionless Keldysh parameter, $\gamma=\frac{m\omega}{|e|E_{0}}$ Keldysh (1965), where $E_{0}$ is the field amplitude, $\omega$ is the field frequency, $e$ is the electron charge, and $m$ is the electron mass. When $\gamma>>1$, the process probability shows a perturbative power-law scaling with field intensity. Instead, for $\gamma<<1$, it exhibits a manifestly non-perturbative exponential dependence on $\frac{1}{E_{0}}$, similar to the case of a constant electric field Nikishov (1969).
Particle production in a spatially homogeneous single pulse of an electric field has been explored Narozhnyi and Nikishov (1970); Nikishov (2001), and methods for tackling particle creation in an arbitrary time-dependent electric field have been developed Narozhnyi and Nikishov (1970); Brezin and Itzykson (1970); Bagrov et al. (1975). Because a single pulse of an electric field is an idealized form of an electric field created by two colliding laser beams, particle generation in an alternating electric field has been explored for a more realistic scenario Mostepanenko and Frolov (1974); Narozhnyi and Nikishov (1973). However, it was found that the mass of the created particle, via the Schwinger mechanism, exponentially suppresses the pair generation rate, necessitating a very strong electric field to observe this phenomenon. Therefore, this event has not yet been the subject of any experimental observations. This makes it unclear how well the theoretical prediction captures the physics of pair production.
Nevertheless, the study of pair production in strong electric fields subject has attracted sustained interest from theoreticians in recent years due to the extraordinary progress in the development of the ultra-intense lasers technique and the strong field QED experimental studies that are planned at upcoming high-intensity laser facilities, such as the European X-Ray Free-Electron Laser Andruszkow et al. (2000), the Extreme-Light Infrastructure Heinzl and Ilderton (2009); Dunne (2009), the Exawatt Center for Extreme Light Studies Khazanov and et al. (2023). The electric field strength is fast getting closer to the critical value $E_{\text{c}}$.
Additionally, it is suggested that the Schwinger mechanism can be indirectly tested in the condensed matter system of a single monolayer of graphene, where the electrons are roughly characterized by the massless pseudo-relativistic Dirac equation Allor et al. (2008); Klimchitskaya and Mostepanenko (2013); Fillion-Gourdeau and MacLean (2015).
Particle production can be viewed as evolving a quantum system from an initial equilibrium configuration to a new final equilibrium configuration via an intermediate non-equilibrium evolution caused by a strong field background. In intermediate states, when matter fields interact with a time-dependent external field, the classical Hamiltonian loses its time-translation invariance, leading to the evolution of the vacuum state into a state where mixing between positive and negative energy states occurs. Special care is needed in identifying “positive” and “negative” energy states in the presence of an external electric field. Thus, energy is no longer a good quantum number, and there is no well-defined quantum number to characterize a “particle”. However, the standard way of describing particle creation by an external field involves the asymptotic analysis of particle states in the remote past (or in-states, before the external field is switched on) and the distant future (or out-states, long after the interaction with the external field has ended). This method is well understood using the quantum field operator expression in terms of creation and annihilation operators, which are connected to single-particle states in both the past and the future. Then, by relating the two sets of operators in the past and future, one can derive an equation for the S-matrix of the process. This equation, in turn, determines the number of particles produced in the process.
Various methodologies have been developed to investigate pair production in the presence of strong fields. These include the proper-time method Schwinger (1951); DeWitt (1975), the canonical method Mottola (1985); Tomaras et al. (2001, 2000), Green’s function techniques Gavrilov and Gitman (1996), semiclassical tunneling Parikh and Wilczek (2000), the Schrödinger-functional approach Hallin and Liljenberg (1995), functional techniques Fried and Woodard (2002); Avan et al. (2003); Ilderton (2014),the Wigner formalismZhuang and Heinz (1996); Hebenstreit et al. (2011a, b),mean-field treatments Smolyansky et al. (1998), and worldline instanton methods Affleck et al. (1982); Kim and Page (2002). Studies on pair production in intense laser fields Alkofer et al. (2001); Roberts et al. (2002) have employed analytical and numerical approaches to predict pair-production rates. However, these investigations primarily focus on pair formation at asymptotic times, which alone is insufficient for fully capturing the dynamical aspects of pair production and its time-dependent effects. Several authors have derived expressions for the number of pairs created after a time $T$, where $T$ exceeds the duration of the electric field pulse or corresponds to a moment after the field has vanished Gavrilov and Gitman (1996); Di Piazza (2004); Melike et al. (2012); Ilderton (2022); Adorno et al. (2017). These expressions provide a reliable approximation of the actual pair count.
@@
Understanding pair production dynamics from the initial to final states requires a framework that accounts for transient and non-equilibrium effects. In this context, the distribution function of quasiparticles, governed by the quantum kinetic equation (QKE), offers an effective description of the time-dependent evolution of pair production Blaschke et al. (2015); Smolyansky et al. (1998). Defining the particle distribution function in the quasiparticle basis is both physically meaningful and intuitive. This approach is grounded in the straightforward physics of how a time-dependent field distorts the time-invariance symmetry of the Hamiltonian, leading to mixing between momentum states of positive and negative energies. This method is particularly relevant for exploring the complex dynamical behavior of pair production in spatially uniform, time-dependent electric fields. Specifically, it has been demonstrated that the temporal evolution of the quasiparticle distribution function in momentum space provides insights into the asymptotic states of the quantum field and offers a comprehensive description of the process during the interaction between the matter field and the strong field background Blaschke et al. (2017). It was shown that the distribution function evolves through three distinct temporal stages: the quasi-electron-positron plasma (QEPP) stage, the transient stage, and finally the residual electron-positron plasma (REPP) stage Banerjee and Singh (2022, 2019). In examining the pair production process, both the QEPP stage, where quasi-particles are formed, and the REPP stage, where real particles emerge, are crucial Blaschke:2013ip; Shen:2018lbq. While the Schwinger mechanism is primarily known for producing elementary particles in a vacuum under external fields, it can also generate quasiparticle excitations in exotic materials such as graphene Allor et al. (2008); Zubkov (2012); Gavrilov et al. (2012); Klimchitskaya and Mostepanenko (2013); Fillion-Gourdeau and MacLean (2015), which are experimentally more accessible, and in superconductors Solinas et al. (2021). Moreover, while ultracold atoms in optical lattices have been investigated as simulators for the Schwinger mechanism Kasper et al. (2016); Halimeh et al. (2025).
In the present work, we consider the production of electron-positron pairs from the vacuum in a time-varying, spatially uniform pulsed electric field given by $E\left(t\right)=E_{0}\textrm{sech}^{2}\left(\frac{t}{\tau}\right)$, with amplitude $E_{0}$ and pulse duration $\tau$. Such background field has received extensive attention in the literature Narozhnyi and Nikishov (1970); Dunne and Hall (1998); Balantekin et al. (1991); Gavrilov and Gitman (1996); Nikishov (2004); Levai and Skokov (2010); Klimchitskaya and Mostepanenko (2013), with a focus on the asymptotic behavior of the probability of pair production. To this end, we study the evolution of the single-particle distribution function, $f(\bm{p},t)$, which is rigorously derived by canonical quantization of the Dirac field and subsequent Bogoliubov transformation to a quasi-particle representation Smolyansky et al. (2017a).
@@
The distribution function exhibits three distinct dynamical stagesBanerjee and Singh (2022); Otto et al. (2016); Nousch et al. (2017); Smolyansky et al. (2017a). By analyzing the temporal evolution of $f(\bm{p},t)$, we find that the emergence and characteristics of these dynamical stages, QEPP, transient, and REPP, are qualitatively and quantitatively influenced by both longitudinal and transverse components of momentum. Next, we analyze the time evolution of the longitudinal momentum spectrum (LMS) of created particles. In the tunneling regime ($\gamma<1$), we observe oscillatory behavior in the LMS at time $t>2\tau$, and this oscillation pattern continuously changes from $t>2\tau$ up to $t<6\tau$. This oscillatory behavior at finite times clearly illustrates the quantum interference effects arising from various dynamical processes and channels contributing to the creation of (quasi-)particles with a given momentum. It can also be understood in terms of the vacuum polarization function $u(\bm{p},t)$ and its counterpart, the depolarization function $v(\bm{p},t)$, which encapsulate the underlying particle production dynamics. We emphasize that the oscillations seen in the LMS are not artifacts but rather possess significant physical relevance. In the multi-photon regime, we observe that LMS at finite times near $t=3\tau$ exhibits a multi-modal structure.
Furthermore, we develop an analytical framework valid for finite times $t>\tau$. Specifically, we derive an analytical expression for the single-particle distribution function as a power series in the small parameter $(1-y)$, where $y=\frac{1}{2}(1+\tanh{(\frac{t}{\tau})})$. The intriguing dynamical features of the momentum distribution function at finite times arise from a function appearing in this expansion. The finite-time behavior of the momentum spectra can be understood in terms of functions that reveal three distinct behaviors, which are discussed in detail in the results section.
Additionally, we explored how the transverse momentum affects LMS and discovered that the oscillatory behavior depends on the value of transverse momentum at that time. Subsequently, we analyzed the temporal progression of the transverse momentum spectrum(TMS) of generated pairs, providing valuable insights into the pair production process. At finite times, the transverse momentum spectra reach their peak at a nonzero value, reflecting the influence of multiphoton transitions in quasiparticle creation.
@@
The preliminary results of this work were presented in Refs. Sah and Singh (2023, 2024). After the report in Ref. Sah and Singh (2023) appeared, one of the authors of Ref. Diez et al. (2023) (C.Kohlfürst, private communication) informed us about their findings, which exhibit some overlap with the results presented here. In this manuscript, we compare and discuss our findings with those of Ref. Diez et al. (2023) whenever applicable.
@@
Throughout the paper, we use natural units and set $\hslash=c=m=|e|=1$, and express all variables in terms of the electron mass unit.
@@ II Theory
In this section, we review the canonical quantization method within the context of a time-dependent electric field, as discussed in Grib et al. (1972). This approach offers valuable insights into pair production phenomena, including detailed information on particle number and momentum distribution at each moment. To begin, we establish solutions for the single-particle scenario governed by the Dirac equation in the presence of such a field. We start by writing the Dirac equation for a particle in an electromagnetic field, which takes the following form:
@@
where $A^{\mu}$ is the four-vector potential, $\Psi(\bm{x},t)$ is the four-component spinor. For Dirac gamma matrices $\gamma^{\mu}$ we chose the Weyl basis Ryder (1996)
$\gamma^{0}=\begin{pmatrix}\mathbb{I}&0\\ 0&-\mathbb{I}\end{pmatrix},\quad\gamma^{i}=\begin{pmatrix}0&-\sigma^{i}\\ \sigma^{i}&0\end{pmatrix},\quad(i=1,2,3),$ (2)
where $\mathbb{I}$ is the identity matrix and $\sigma^{i}$ are the Pauli matrices. The Dirac gamma matrices $\gamma^{\mu}$ satisfy the anti-commutation relations:
@@
$\left[(i\partial_{\mu}-eA_{\mu})^{2}-\frac{e}{2}\sigma^{\mu\nu}\mathcal{F}_{\mu\nu}-m^{2}\right]\chi(\bm{x},t)=0,$ (6)
where $\mathcal{F}_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$ is the field strength tensor, $\sigma^{\mu\nu}=\frac{i}{2}[\gamma^{\mu},\gamma^{\nu}]$ is the commutator of the gamma matrices, and $\chi(\bm{x},t)$ is a four component spinor.
@@
Spatial homogeneity implies that the solutions of the form
@@
Subsequently, Eq. (7) reduces to
$\left(\partial_{t}^{2}+\textrm{i}eE(t)\gamma^{0}\gamma^{3}+\omega^{2}(\bm{p},t)\right)\chi_{\bm{p}}(t)=0,$ (9)
where the time-dependent frequency is
$\omega(\bm{p},t)=\sqrt{\epsilon_{\perp}^{2}(p_{\perp})+P^{2}(p_{\parallel},t)}.$ (10)
Here, the transverse energy is given by $\epsilon_{\perp}(p_{\perp})=\sqrt{m^{2}+p_{\perp}^{2}}$, and the longitudinal kinetic momentum is $P(p_{\parallel},t)=p_{\parallel}-eA(t)$. The longitudinal momentum component $p_{\parallel}\equiv p_{3}$ is directed along the external electric field, while the transverse momentum magnitude $p_{\perp}=\sqrt{p_{1}^{2}+p_{2}^{2}}$ corresponds to the vector $\bm{p}_{\perp}$, which lies in the plane perpendicular to the field.
@@
These solutions will be interpreted as describing an electron ($\lambda=+$), and its antiparticle, i.e., a positron ($\lambda=-$).
@@
where the spinor solutions, $\Psi_{{\bm{p}}r}^{(\lambda)}(\bm{x},t)$ are normalized according to the product:
@@
Now, the Hamiltonian can be calculated from the energy-momentum tensor, which yields
@@
To calculate the spectrum, the Hamiltonian is diagonalized through a basis transformation to the quasi-particle representation using the new time-dependent operators $\hat{B}_{{\bm{p}}r}(t)$ and $\hat{D}_{{\bm{p}}r}(t)$. The relation between the $\hat{b}_{{\bm{p}}r},\hat{d}_{{\bm{p}}r}$and $\hat{B}_{{\bm{p}}r}(t),\hat{D}_{{\bm{p}}r}(t)$ operators is given by the Bogoliubov transformation Grib et al. (1972); Schmidt et al. (1998)
@@
$\hat{D}_{{\bm{p}}r}(t)$ $=\alpha_{-{\bm{p}}}(t)\hat{d}_{{\bm{p}}r}-\beta_{-{\bm{p}}}(t)\hat{b}_{-{\bm{p}}r}^{\dagger},$ (29)
@@
Here, the operators $\hat{B}_{{\bm{p}}r}(t)$ and $\hat{D}_{{\bm{p}}r}(t)$ describe the creation and annihilation of quasiparticles at time $t$ with respect to the instantaneous vacuum state Schmidt et al. (1998). The substitution of Eqs. (28) and (29) into Eq. (21) results in a new representation of the field operator
@@
$\phi_{\bm{p}}^{(\lambda)}(t)=\frac{e^{-\textrm{i}\lambda\Theta_{\bm{p}}(t_{0},t)}}{\sqrt{2\omega(\bm{p},t)\bigl{(}\omega(\bm{p},t)-\lambda P(p_{\parallel},t)\bigr{)}}},$ (35)
with the dynamical phase
@@
The lower limit $t_{0}$ of the integral in Eq. (36) is not fully determined as it only fixes an arbitrary phase. Note that the functions $\phi_{\bm{p}}^{(\pm)}(t)$ are chosen such that they coincide with the mode functions $\psi_{\bm{p}}^{(\pm)}(t)$ in the case of a vanishing electric field Smolyansky et al. (1998); Kluger:1998bm.
@@
We now define the occupation number of electrons in the given eigenmode $\bm{p}r$ of the fermionic field using the time-dependent creation and annihilation operators for the initial vacuum state as
$f_{r}(\bm{p},t)$ $=\langle{0_{in}}|\hat{B}_{{\bm{p}}r}^{\dagger}(t)\hat{B}_{{\bm{p}}r}(t)|{0_{in}}\rangle,$ (41)
@@
$\bar{f}_{r}(-\bm{p},t)$ $=\langle{0_{in}}|\hat{D}_{-{\bm{p}}r}^{\dagger}(t)\hat{D}_{-{\bm{p}}r}(t)|{0_{in}}\rangle.$ (42)